OpenStudy (anonymous):

What is the solution to the rational equation (x)/(2x-1) + 1/4 = 2/2x-1?

OpenStudy (anonymous):

\[\frac{ x }{ 2x-1 } + \frac{ 1 }{ 4 } = \frac{ 2 }{ 2x-1 }\]

OpenStudy (marihelenh):

First, you need to get a common denominator.

OpenStudy (anonymous):

My options are here:

OpenStudy (anonymous):

ok so you'd multiply everything by 2x-1 right?

OpenStudy (marihelenh):

Yes, but also by 4, because that is a part of the denominator and it needs to equal the rest.

OpenStudy (anonymous):

ok so then what would that look like

OpenStudy (tkhunny):

FIRST, you write down \(x \ne \dfrac{1}{2}\).

OpenStudy (anonymous):

like after you multiply everything by 2x-1, you multiply all that by 4?

OpenStudy (marihelenh):

\[(\frac{ x }{ 2x-1 })(\frac{ 4 }{ 4 })\] \[(\frac{ 1 }{ 4 })(\frac{ 2x-1 }{ 2x-1 })\] \[(\frac{ 2x }{ 2x-1 })(\frac{ 4 }{ 4 })\]

OpenStudy (marihelenh):

You need to solve these out, that way each has the same denominator.

OpenStudy (anonymous):

ohhh ok. so, i did mental math real quick and i thiiiink the answer is 9/2?

OpenStudy (anonymous):

or maybe 3/2. its kinda late so im tryna do this fast lol

OpenStudy (anonymous):

um is that correct? or should i try again

OpenStudy (marihelenh):

Just fill in your answers for x and check.

OpenStudy (anonymous):

kk ill finish in a sec

OpenStudy (anonymous):

ok im p sure it's 3/2

OpenStudy (marihelenh):

Yep! Great job!

OpenStudy (anonymous):

ok great thanks